# External charged debris in a flowing plasma : charge fluctuation induced complexity

### v1

@@ Abstract
organization=Affiliation: Physics Department, Gauhati University,addressline=Gauhati University, city=Guwahati, postcode=781014, state=Assam, country=India
@@ 1 Introduction
The paper is organized as follows. In Section II, we present the results of kinetic and fluid simulations with the help of our h-PIC-MMC and mFCT codes, respectively. In Section III, we provide a theoretical analysis of the results obtained in Section II. We provide an additional analysis of the results obtained in earlier sections, in terms of power spectrum analysis in Section IV. In Section V, we conclude. We also provide a brief analysis of the effect of fluctuation of debris charge on turbulent dynamics.
@@ 2.1 The premise
Though technically it should be possible to model a self-consistent periodic fluctuation of charge in an external debris embedded in a plasma, in reality it is much harder to model it mathematically as it would naturally require calculation of collision cross-sections to exactly determine the electron and ion currents to the debris, which may not obey the much used orbit motion limited (OML) theory, usually used to calculate the currents. It may be further complicated by extended and arbitrary shape of the debris. Besides, a self-consistent fluctuation will necessarily require a coupling a circuit-like differential equations with the partial differential equations of the fluid model for the plasma. So, for a fluid simulation, we shall require a charging equation like
@@ 2.2 The plasma model
2.3 Kinetic model - PIC simulation
@@ 2.5 Nonlinear Landau damping?
Figure 7: Flattening of the ion VDF due to fluctuation. While the top panels (a) and (b) display respectively the VDFs calculated from the PIC simulation without $(\nu=0)$ and with $(\nu\sim 1)$ fluctuation, the bottom panels (c) and (d) show the numerically calculated FWHMs for the corresponding VDFs. The panel (d) on the right clearly shows that the VDF is flattening due to fluctuation.

### v2

@@ Abstract
organization=Physics Department, Gauhati University,city=Guwahati, postcode=781014, state=Assam, country=India
{highlights}
Model: Investigates the effect of periodic charge fluctuation of external debris, induced by ion-acoustic wake oscillations, in a flowing $e$-$i$ plasma.
Nonlinear complexity: Demonstrates emergence of debris charge-dependent chaos and nonlinear Landau damping (NLLD) using kinetic and fluid simulations.
Nonlinear modelling: Theoretical analysis confirms simulation findings.
New insights: Positively charged debris induces stronger localised chaos compared to negatively charged debris, with latter exhibiting NLLD.
Potential application: These new insights into plasma-debris interactions may be useful for management of space debris in low Earth orbit.
@@ 1 Introduction
The paper is organized as follows. In Section II, we present the results of kinetic and fluid simulations with the help of our h-PIC-MCC and mFCT codes, respectively. In Section III, we provide a theoretical analysis of the results obtained in Section II. We provide an additional analysis of the results obtained in earlier sections, in terms of power spectrum analysis in Section IV. In Section V, we conclude. We also provide a brief analysis of the effect of fluctuation of debris charge on turbulent dynamics.
@@ 2.1 The premise
Though technically it should be possible to model a self-consistent periodic fluctuation of charge in an external debris embedded in a plasma, in reality it is much harder to model it mathematically as it would naturally require calculation of collision cross-sections to exactly determine the electron and ion currents to the debris, which may not obey the much used orbit motion limited (OML) theory, usually used to calculate the currents. It may be further complicated by extended and arbitrary shape of the debris. Besides, a self-consistent fluctuation will necessarily require a coupling of a circuit-like differential equation with the partial differential equations of the fluid model for the plasma, so that we shall require a charging equation like
@@ 2.2 The plasma model
2.3 Kinetic model – PIC simulation
@@ 2.5 Nonlinear Landau damping?
Figure 7: Flattening of the ion VDF due to fluctuation. While the top panels (a) and (b) display respectively the VDFs calculated from the PIC simulation without $(\nu=0)$ and with $(\nu\sim 0.7)$ fluctuation, the bottom panels (c) and (d) show the numerically calculated FWHMs for the corresponding VDFs. The panel (d) on the right clearly shows that the VDF is flattening due to fluctuation.
@@ 5 Conclusions
Acknowledgement
BJ would like to acknowledge Gauhati University for research scholarship grant GU/UGC/GURF/2023/712. HS would like to thank CSIR-HRDG, New Delhi, India for Senior Research Fellowship research grant 09/059(0074)/2021-EMR-I.